OpenStudy (anonymous):

similar to the last. from $48.50 to $38.80

OpenStudy (anonymous):

38.80 = 100% 48.50 - 38.80 = 9.7 9.7/38.80 = 0.25 = 25% 38.8 + 9.7 = 48.50 100% + 25% = 125% 48.50 = 125% of 38.80